# Scaling limit of heavy-tailed nearly unstable cumulative INAR($\infty$) processes and rough fractional diffusions

### v6

@@ 2.1 Heuristic Motivation for the Scaling Limit
$\gamma_{T}=\frac{1}{\sqrt{\mu^{T}T(1-a_{T})}},$
then (16) can be written as
$\frac{1}{\left\lVert A^{T}\right\rVert_{1}\mu^{T}}\kappa_{\lfloor sT\rfloor}^{T}\approx$ $\gamma_{T}v_{T}\int_{0}^{s}\left(s-u\right)^{\alpha-1}E_{\alpha,\alpha}\left(-v_{T}\left(s-u\right)^{\alpha}\right)\sqrt{C_{u}^{T}}dB_{u}^{T}$ (17)
$\stackrel{{\scriptstyle\text{w}}}{{\rightarrow}}$ $\gamma_{\infty}v_{\infty}\int_{0}^{s}(s-u)^{\alpha-1}E_{\alpha,\alpha}(-v_{\infty}(s-u)^{\alpha})\sqrt{C_{u}^{\infty}}dB_{u},$
as $T\rightarrow\infty$, where
$\gamma_{\infty}=\frac{1}{\sqrt{\lambda\nu^{*}}},v_{\infty}=\lambda.$
2.2 A Microstructural Foundation for the Initial Value
In practical applications, particularly for financial models like the rough Heston model, specifying the initial value of the volatility process, $\dot{Y}_{0}=v_{0}$, is essential. This is fundamentally different from setting the initial value of the integrated process, $Y_{0}$. Achieving this requires embedding the initial dynamics directly into the discrete-time microstructural model.
Following the methodology of El Euch and Rosenbaum (2019) and its discrete-time adaptation in Wang and Cui (2025), we introduce a carefully constructed time-dependent baseline intensity, $\hat{\mu}_{T}(n)$. This baseline is no longer a constant but is designed to create a burst of initial activity that, upon scaling, converges to the desired initial value $v_{0}$.
Let us redefine the intensity process from (3) in Section 2. The immigration component is now time-varying:
$\lambda_{n}^{T}=\hat{\mu}_{T}(n)+\sum_{s=1}^{n-1}\eta_{n-s}^{T}X_{s}^{T},\quad n\in\mathbb{N},$ (18)
where the time-dependent baseline intensity is given by:
Definition 1 (Time-Dependent Baseline for Initial Value).
For a given constant $\xi\geq 0$, the baseline intensity is defined as:
$\hat{\mu}_{T}(n):=\mu^{T}+\xi\mu^{T}\left(\frac{1}{1-a_{T}}\left(1-\sum_{s=1}^{n-1}\eta_{s}\right)-\sum_{s=1}^{n-1}\eta_{s}\right),$
where $(\eta_{n})_{n\geq 1}$ is the base kernel sequence (before scaling by $a_{T}$) satisfying $\sum_{n=1}^{\infty}\eta_{n}=1$. The parameter $\xi$ is a dimensionless constant that will directly control the initial value of the limit process.
The intuition for the above definition is as follows. At the very beginning, for $n=1$, the sums are empty, leading to a large initial baseline intensity:
$\hat{\mu}_{T}(1)=\mu^{T}\left(1+\frac{\xi}{1-a_{T}}\right).$
Since $(1-a_{T})^{-1}$ is of order $\mathcal{O}(T^{\alpha})$, this creates a significant initial burst of events. As $n$ increases, the term in the parenthesis decays, and $\hat{\mu}_{T}(n)$ quickly converges to the constant baseline $\mu^{T}$. This initial, engineered transient behavior is precisely what generates the non-zero starting value for the rate process $\dot{Y}_{t}$ in the continuous-time limit.
Now Theorem 1 can be extended to Corollary 2 which is given below. The Corollary shows that under this modified baseline intensity, the limit process $\dot{Y}_{t}$ starts from a value directly determined by $\xi$. We first introduce a Lemma which is useful in proving Corollary 2.
Lemma 9 (Renewal Expansion with Time-Dependent Baseline).
Under the time-dependent baseline intensity $\hat{\mu}_{T}(n)$ from Definition 1, the cumulative intensity and the centered cumulative process have the following representations for $n\geq 1$:
$\mathbb{E}[N_{n}^{T}]=\sum_{i=1}^{n}\left(\hat{\mu}_{T}(i)+\sum_{s=1}^{i-1}A_{i-s}^{T}\hat{\mu}_{T}(s)\right),$ (19)
and
$N_{n}^{T}-\mathbb{E}[N_{n}^{T}]=M_{n}^{T}+\sum_{s=1}^{n-1}A_{n-s}^{T}M_{s}^{T}.$ (20)
@@ 2.2 A Microstructural Foundation for the Initial Value
We provide the proof in Section 4.8. ∎
Corollary 2 (Characterization of the Derivative Process with Initial Process Value).
Under the assumptions of Theorem 1 and with the baseline intensity from Definition 1, the limit process $(\dot{Y}_{t})_{t\in[0,1]}$ is the unique solution to the stochastic Volterra equation
$\dot{Y}_{t}=F^{\alpha,\lambda}(t)+\xi\left(1-F^{\alpha,\lambda}(t)\right)+\frac{1}{\sqrt{\nu^{*}\lambda}}\int_{0}^{t}f^{\alpha,\lambda}(t-s)\sqrt{\dot{Y}_{s}}dB_{s},$ (21)
@@
We provide the proof in Section 5.3. ∎
This framework provides a theoretically sound and practical method for simulating the fractional CIR process from any desired starting level $v_{0}$ by directly encoding this value into the microscopic dynamics.
@@ 3 Numerical Validation and Application
3.1 Simulation via INAR($\infty$) Approximation
Our main theoretical result, encapsulated in Theorem 1 and Corollary 2, provides a novel and efficient pathway for simulating the fractional CIR process. The core idea is that we can generate sample paths of the limiting continuous-time process by simulating its discrete-time antecedent—the heavy-tailed cumulative INAR($\infty$) process—for a large but finite $T_{\text{steps}}$. This approach leverages the computational simplicity of simulating discrete count processes, which only requires iterative sampling from a Poisson distribution.
A key feature of our method is the ability to specify a non-zero initial value, $v_{0}$, for the fractional CIR process ($\dot{Y}_{t}$). This is not achieved by an ad-hoc adjustment post-simulation, but is instead built directly into the microstructural dynamics of the INAR($\infty$) process. As outlined in Definition 1, we employ a time-dependent baseline intensity, $\hat{\mu}_{T}(n)$, where the parameter $\xi$ is set to the desired initial value $v_{0}$. This creates an initial burst of activity that correctly sets the starting point of the process in the continuous-time limit.
@@ 3.1 Simulation via INAR($\infty$) Approximation
1:Input: Target initial value $v_{0}$, parameters $\alpha$, $\lambda_{\text{param}}$, $\nu^{*}$, and number of time steps $T_{\text{steps}}$.
@@
6:  $\mu_{T}=\nu^{*}\cdot\delta^{-1}\cdot T_{\text{steps}}^{\alpha-1}$
7:  $a_{T}=1-(\lambda_{\text{param}}\cdot\delta)/T_{\text{steps}}^{\alpha}$
@@
11:  Calculate the time-dependent baseline intensity $\hat{\mu}_{T}(n)$ using the formula from Definition 1.
@@
14:  Draw the count for the current step: $X_{n}^{T}\sim\text{Poisson}(\max(\lambda_{n}^{T},0))$.
@@
Figure 1: A sample path of the fractional CIR process $(V_{t}=\dot{Y}_{t})_{0\leq t\leq 1}$, simulated via the INAR($\infty$) approximation. The simulation parameters are $\alpha=0.62$, $\lambda_{\text{param}}=1.0$, $\nu^{*}=200.0$, and a target initial value $v_{0}=0.7$. The number of discrete steps is $T_{\text{steps}}=5000$. The red dashed line indicates the target initial value, which the simulated path closely matches at $t=0$.
@@ 3.2 A Goodness-of-Fit Test and application in rough Heston models
Direct validation of the univariate fractional CIR process is challenging due to a scarcity of reliable benchmarks; therefore, we first highlight the framework’s successful application to its well-benchmarked bivariate counterpart, the rough Heston model. To rigorously test the accuracy and goodness-of-fit of our INAR($\infty$) approximation, we focus on the rough volatility case by setting the parameter to $\alpha=0.62$.
@@
The robustness of the underlying INAR($\infty$) framework is highlighted by its successful application in related literature. In the work of Wang and Cui (2025), for instance, a bivariate INAR($\infty$) process is developed to model the interacting dynamics of buy and sell order flows, providing a microstructural foundation for the full rough Heston model. The authors there prove the model’s convergence and validate its accuracy through extensive numerical experiments. The success of the framework in that more complex, bivariate setting (as shown in Wang and Cui (2025)) provides strong collateral support for the validity of our approach. We therefore proceed with a direct numerical validation, comparing our simulation results against the benchmark from Callegaro et al. (2021).
@@
Table 1: Comparison of European option prices for the rough Heston model ($\alpha=0.62$) between the benchmark solution and our INAR($\infty$) simulation. The simulation uses 100,000 paths with $T=250$.
@@
While we use European options in this test for validation against a well-established benchmark, a key advantage of our simulation framework is its versatility. Being a Monte Carlo-based method, it naturally extends to the pricing of path-dependent options—such as Asian, Lookback, and Barrier options—for which analytical or semi-analytical solutions are often intractable. Moreover, the computational efficiency of this approach is a significant benefit. As demonstrated in the bivariate context by Wang and Cui (2025), the INAR($\infty$) simulation framework is inherently parallelizable and proves to be considerably faster than traditional discretization schemes, such as the Euler method, applied to the limiting stochastic Volterra equation. It is stated in Wang and Cui (2025) that the total execution time on a MacBook M2 Pro is 18.3068 seconds for calculating all four options (European, Asian, Lookback, Barrier) for 100000 paths with $T=250$.
4 Proof of Lemmas
4.1 Proof of Lemma 1

### v7

@@ 2.3 Heuristic Motivation for the Scaling Limit
$N_{n}^{T}-\mathbb{E}[N_{n}^{T}]=M_{n}^{T}+\sum_{s=1}^{n-1}A_{n-s}^{T}M_{s}^{T}.$ (2.16)
@@
We provide the proof in Section A.9. ∎
Corollary 2.18 (Characterization of the Derivative Process with Initial Process Value).
Under the assumptions of Theorem 2.13 and with the baseline intensity from Definition 2.16, the limit process $(\dot{Y}_{t})_{t\in[0,1]}$ is the unique solution to the stochastic Volterra equation
$\dot{Y}_{t}=F^{\alpha,\lambda}(t)+\xi\left(1-F^{\alpha,\lambda}(t)\right)+\frac{1}{\sqrt{\nu^{*}\lambda}}\int_{0}^{t}f^{\alpha,\lambda}(t-s)\sqrt{\dot{Y}_{s}}dB_{s},$ (2.17)
@@
We provide the proof in Section B.3. ∎
@@ 3 Numerical Validation and Application
3.1 Simulation via INAR Approximation
Our main theoretical result, encapsulated in Theorem 2.13 and Corollary 2.18, provides a novel pathway for simulating the fractional CIR process. The core idea is that we can generate sample paths of the limiting continuous-time process by simulating its discrete-time antecedent—the heavy-tailed cumulative INAR($\infty$) process—for a large but finite $T_{\text{steps}}$. This approach leverages the computational simplicity of simulating discrete count processes, which only requires iterative sampling from a Poisson distribution.
A key feature of our method is the ability to specify a non-zero initial value, $v_{0}$, for the fractional CIR process ($\dot{Y}_{t}$). This is not achieved by an ad-hoc adjustment post-simulation, but is instead built directly into the microstructural dynamics of the INAR($\infty$) process. As outlined in Definition 2.16, we employ a time-dependent baseline intensity, $\hat{\mu}_{T}(n)$, where the parameter $\xi$ is set to the desired initial value $v_{0}$. This creates an initial burst of activity that correctly sets the starting point of the process in the continuous-time limit.
@@ 3.1 Simulation via INAR Approximation
1:Input: Target initial value $v_{0}$, parameters $\alpha$, $\lambda$, $\nu^{*}$, and number of time steps $T_{\text{steps}}$.
@@
6:  $\mu^{T}=\nu^{*}\cdot\delta^{-1}\cdot T_{\text{steps}}^{\alpha-1}$
7:  $a_{T}=1-(\lambda\cdot\delta)/T_{\text{steps}}^{\alpha}$
@@
11:  Calculate the time-dependent baseline intensity $\hat{\mu}_{T}(n)$ using the formula from Definition 2.16.
@@
14:  Draw the count for the current step: $X_{n}^{T}\sim\text{Pois}(\max(\lambda_{n}^{T},0))$.
@@
Figure 1: A sample path of the fractional CIR process $(V_{t}=\dot{Y}_{t})_{0\leq t\leq 1}$, simulated via the INAR($\infty$) approximation. The simulation parameters are $\alpha=0.62$, $\lambda=1.0$, $\nu^{*}=200.0$, and a target initial value $v_{0}=0.7$. The number of discrete steps is $T_{\text{steps}}=5000$. The red dashed line indicates the target initial value, which the simulated path closely matches at $t=0$.
@@ 3.2 A Goodness-of-Fit Test and application in rough Heston models
Direct validation of the univariate fractional CIR process is challenging due to a scarcity of reliable benchmarks; therefore, we first highlight the framework’s successful application to its well-benchmarked bivariate counterpart, the rough Heston model. Recall that under the risk-neutral measure, the rough Heston model specifies the asset price $S$ and the instantaneous variance $V$ by
$dS_{t}=S_{t}\sqrt{V_{t}}\,dW_{t},\qquad V_{t}=V_{0}+\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}\gamma\left(\theta-V_{s}\right)\,ds+\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}\gamma\nu\sqrt{V_{s}}\,dB_{s},$
where $\alpha\in(1/2,1)$ controls the roughness (Volterra memory), $\gamma>0$ is the mean-reversion speed, $\theta>0$ the long-run variance level, $\nu>0$ the volatility-of-volatility, and $\rho\in(-1,1)$ is the correlation between the Brownian motions $(W,B)$. For consistency with the benchmark parametrization used later in the numerical section, we employ the standard reparametrization in terms of $(\beta,\mu,\xi)$:
$\rho=\frac{1-\beta}{\sqrt{2(1+\beta^{2})}}\quad\Longleftrightarrow\quad(2\rho^{2}-1)\beta^{2}+2\beta+(2\rho^{2}-1)=0,$
hence
$\beta=\frac{1\pm 2|\rho|\sqrt{1-\rho^{2}}}{1-2\rho^{2}}.$
$\mu=\frac{\theta(1+\beta^{2})}{\gamma\nu^{2}(1+\beta)^{2}},\qquad\xi=\frac{V_{0}}{\theta}.$
so that the correlation/leverage and the variance-level inputs $(\rho,\theta,V_{0})$ are matched exactly by the derived parameters $(\beta,\mu,\xi)$ reported below. To rigorously test the accuracy and goodness-of-fit of our INAR($\infty$) approximation, we focus on the rough volatility case by setting the parameter to $\alpha=0.62$.
We emphasize that our main theoretical results concern a univariate limit. However, for the univariate fractional CIR/Volterra variance dynamics there is, to the best of our knowledge, no closed-form or widely used high-accuracy benchmark that would allow a quantitative goodness-of-fit assessment comparable to the option-pricing benchmarks available for rough Heston. We therefore validate the INAR($\infty$) simulation methodology in the well-benchmarked rough Heston setting, where accurate reference prices are available (e.g. Callegaro et al. (2021)). This numerical experiment is included as a consistency check for the microstructural INAR($\infty$) approximation and its simulation accuracy.
  • Simulation setup. We use the standard rough Heston parameter set from Callegaro et al. (2021) (also used in Wang et al. (2025)): $\gamma=0.1$, $\rho=-0.681$, $\nu=0.331$, $\theta=0.3156$, $V_{0}=0.0392$, $S_{0}=100$, and the corresponding derived parameters $\beta=27.5583$, $\mu=26.8592$, and $\xi=0.124208$. We simulate with $T=320$ time steps per year (i.e. $\Delta t=1/T$).
  • Bivariate microstructural setting (Table 1). To simulate the rough Heston benchmark, we use the bivariate INAR($\infty$) microstructural model introduced in Wang et al. (2025). In that model, the reproduction coefficients are given by a matrix-valued kernel $(\eta_{n}^{T})_{n\geq 1}$ with $\eta_{n}^{T}\in\mathbb{R}_{+}^{2\times 2}$, defined as $\eta_{n}^{T}=a_{T}\eta_{n}$. The baseline kernel $(\eta_{n})_{n\geq 1}$ is heavy-tailed (entrywise) and normalized so that the associated reproduction operator is critical (spectral radius equal to $1$), while $a_{T}\uparrow 1$ enforces near-instability as in Assumptions 2.12.3. We refer to Wang et al. (2025) for the full construction linking the two-dimensional order-flow process to the rough Heston variance/price dynamics; here we only use this setting as a numerical validation benchmark for the INAR($\infty$) simulation methodology. This bivariate experiment does not affect the univariate limit theory developed earlier.
  • Confidence intervals. For each strike, the reported INAR($\infty$) price is the Monte Carlo mean of the (discounted) payoff over $M$ simulated paths, and the 95% confidence interval is computed as
    $\widehat{P}\ \pm\ 1.96\,\frac{\widehat{\sigma}}{\sqrt{M}},$
    where $\widehat{\sigma}^{2}$ is the sample variance of the simulated discounted payoffs.
@@
The robustness of the underlying INAR($\infty$) framework is highlighted by its successful application in related literature. In the work of Wang et al. (2025), for instance, a bivariate INAR($\infty$) process is developed to model the interacting dynamics of buy and sell order flows, providing a microstructural foundation for the full rough Heston model. The authors there prove the model’s convergence and validate its accuracy through extensive numerical experiments. The success of the framework in that more complex, bivariate setting (as shown in Wang et al. (2025)) provides strong collateral support for the validity of our approach. We therefore proceed with a direct numerical validation, comparing our simulation results against the benchmark from Callegaro et al. (2021).
@@
Table 1: Comparison of European option prices for the rough Heston model ($\alpha=0.62$) between the benchmark solution and our INAR($\infty$) simulation. The simulation uses 500,000 paths with $T=320$.
@@
While we use European options in this test for validation against a well-established benchmark, a key advantage of our simulation framework is its versatility. Being a Monte Carlo-based method, it naturally extends to the pricing of path-dependent options—such as Asian, Lookback, and Barrier options—for which analytical or semi-analytical solutions are often intractable. Moreover, the computational efficiency of this approach is a significant benefit. As demonstrated in the bivariate context by Wang et al. (2025), the INAR($\infty$) simulation framework is inherently parallelizable and proves to be considerably faster than traditional discretization schemes, such as the Euler method, applied to the limiting stochastic Volterra equation. It is stated in Wang et al. (2025) that the total execution time on a Mac mini M4 is 70.3283 seconds for calculating all four options (European, Asian, Lookback, Barrier) for 500,000 paths with $T=320$.
4 Future Work
Acknowledgement
Declarations
  • Author contribution: Yingli Wang led the research, authored the main manuscript, provided the majority of the proofs, prepared Figure 1, and developed Algorithm 1. Qinghua Wang contributed to the proof of Lemma 2.6 and played a key role in restructuring the manuscript. Ping He and Chunhao Cai supervised the project, provided critical revisions, and ensured the overall quality and coherence of the manuscript.
Funding Declarations
Appendix A Proof of Lemmas
A.1 Proof of Lemma 2.2
The proof consists of two main steps. First, we rigorously show that for a function $f(t):=t^{\beta}l(t)$, which is regularly varying with index $\beta$}, its (partial) sum is asymptotically equivalent to the corresponding integral over unit intervals. Second, we apply Karamata’s Integral Theorem(s).
Case (i): $\beta>-1$. First, let’s prove that $\sum_{t=X}^{x}f(t)\sim\int_{X}^{x}f(t)\,dt$. Our goal is to prove that
$\lim_{x\to\infty}\frac{\sum_{t=X}^{x}f(t)}{\int_{X}^{x}f(t)\,dt}=1.$
We use the Stolz–Cesàro theorem. Let $a_{x}=\sum_{t=X}^{x}f(t)$ and $b_{x}=\int_{X}^{x}f(t)\,dt$. Since $l$ is positive, $f(t)=t^{\beta}l(t)$ is positive for all $t\geq X$. Hence $b_{x}$ is strictly increasing, and since $\beta>-1$ we have $b_{x}\to\infty$ as $x\to\infty$, satisfying the conditions for the theorem. We therefore analyze the limit of the ratio of differences:
According to the Uniform Convergence Theorem for regularly varying functions (Bingham et al., 1989, Theorem 1.5.2), the convergence of $f(\lambda x)/f(x)\to\lambda^{\beta}$ is uniform for $\lambda$ in any compact interval $[a,b]$ with $0<a\leq b<\infty$. Let us choose such an interval that contains $1$, for example $[1/2,2]$. For $x$ large enough, our interval [11/x,1][1-1/x,1] is contained within $[1/2,2]$.
$(1-\varepsilon)f(x)\leq f(t)\leq(1+\varepsilon)f(x).$
This demonstrates that $\int_{x-1}^{x}f(t)\,dt\sim f(x)$ as $x\to\infty$. Therefore, the limit of the ratio of differences is 1. By the Stolz–Cesàro theorem, the limit of the original ratio is also 1, which proves that $\sum_{t=X}^{x}f(t)\sim\int_{X}^{x}f(t)\,dt$.
Next, we apply Karamata’s Integral Theorem. Karamata’s theorem for integrals of regularly varying functions (Bingham et al., 1989, Proposition 1.5.8) states that for ${\beta}>-1$,
$\int_{X}^{x}t^{\beta}l(t)\,dt\sim\frac{1}{{\beta}+1}x^{{\beta}+1}l(x)\quad\text{as }x\to\infty.$
Combining the above asymptotic equivalences, we obtain
$\sum_{t=X}^{x}t^{\beta}l(t)\sim\frac{1}{{\beta}+1}x^{{\beta}+1}l(x),$
which proves item (i).
Case (ii): $\beta<-1$. In this case, $f(t)=t^{\beta}l(t)$ is eventually positive and integrable at infinity. We first show that
$\sum_{t=x+1}^{\infty}f(t)\sim\int_{x}^{\infty}f(t)\,dt.$}
Define $A_{x}:=\sum_{t=x+1}^{\infty}f(t)$ and $B_{x}:=\int_{x}^{\infty}f(t)\,dt$. Then $A_{x}\downarrow 0$ and $B_{x}\downarrow 0$ as $x\to\infty$. We again use the Stolz–Cesáro theorem applied to the decreasing sequences $(A_{x})$ and $(B_{x})$, and consider
$\lim_{x\to\infty}\frac{A_{x}-A_{x+1}}{B_{x}-B_{x+1}}=\lim_{x\to\infty}\frac{f(x+1)}{\int_{x}^{x+1}f(t)\,dt}.$}
As in Case (i), by writing $t=\lambda(x+1)$ for $t\in[x,x+1]$ and using the Uniform Convergence Theorem (Bingham et al., 1989, Theorem 1.5.2), we obtain $\int_{x}^{x+1}f(t)\,dt\sim f(x+1)$. Hence the above limit equals $1$, and Stolz–Cesáro yields $A_{x}/B_{x}\to 1$, i.e. $\sum_{t=x+1}^{\infty}f(t)\sim\int_{x}^{\infty}f(t)\,dt$.
Next, we apply Karamata’s theorem for tails of integrals of regularly varying functions (Bingham et al., 1989, Proposition 1.5.10), which states that for $\beta<-1$,
[…]
Combining the last two asymptotic relations proves item (ii):
$\sum_{t=x+1}^{\infty}t^{\beta}l(t)\sim-\frac{1}{\beta+1}\,x^{\beta+1}l(x).$}
This completes the proof.
A.2 Proof of Lemma 2.4
First, note that $\rho^{T}$ is a non-negative step function. Since the renewal (resolvent) sequence was originally defined only for indices $n\geq 1$ by
$A_{n}^{T}:=\sum_{m=1}^{\infty}(\eta^{T*m})_{n},\qquad n\geq 1,$
we now extend it to $n=0$ by the convention
$\eta_{0}:=0,\qquad A_{0}^{T}:=0,$
solely to write the step-function formula for $\rho^{T}$ uniformly over all intervals $[k/T,(k+1)/T)$ with $k\geq 0$. With this convention, $\rho^{T}$ is constant on $[k/T,(k+1)/T)$ with value $TA_{k}^{T}/\|A^{T}\|_{1}$, where $\|A^{T}\|_{1}:=\sum_{n\geq 1}A_{n}^{T}$. Therefore,
$\int_{0}^{\infty}\rho^{T}(x)\,dx$ $=\sum_{k=0}^{\infty}\int_{k/T}^{(k+1)/T}T\frac{A_{k}^{T}}{\|A^{T}\|_{1}}\,dx=\sum_{k=0}^{\infty}\frac{A_{k}^{T}}{\|A^{T}\|_{1}}$
$=\frac{A_{0}^{T}+\sum_{k\geq 1}A_{k}^{T}}{\|A^{T}\|_{1}}=\frac{0+\|A^{T}\|_{1}}{\|A^{T}\|_{1}}=1,$
so $\rho^{T}$ is indeed a probability density on $0,\infty)$.